nathan drew a circle on a grid with a radius of 4 units and the points 2,-1 as its center. What are the pairs of coordinates represented points on the circle
Answer by stanbon(75887)

nathan drew a circle on a grid with a radius of 4 units and the points 2,-1 as its center. What are the pairs of coordinates represented points on the circle
-------------------------------
Equation:
(x-2)^2 + (y+1)^2 = 16
------
Find the x and the y intercept:
-------
Let x = 0,
then 4 + (y+1)^2 = 16
(y+1)^2= 12
y+1 = +-2sqrt(3)
y = -1+-2sqrt(3)
=======================
Let y = 0
then (x-2)^2 + 1 = 16
(x-2)^2 = 15
x-2 = +-sqrt(15)
x = 2+-sqrt(15)
-----------------------
Cheers,
Stan H.